A bag contains 10 red chips and 6 blue chips. Two chips are selected randomly without replacement from the bag.

a) Find the probability that the second is a red chip, given that the first was a blue chip.

b) Find the probability that the two chips have different colors.
Answer #1

a) There are 10 red chips and 6 blue chips

We asked P( 2 Nd red / 1st blue) = (6/16)*( 10/15)

P( 2nd red/1st blue) = 0.25

b) probability of two chips have different colors

P( two have different colors) =( 6/16)*(10/15) + (10/16)*(6/15)

P( two have different colors) = 0.25 + 0.25 = 0.50